OpenStudy (s00411855):

how do you factor 5y^2-125???

OpenStudy (anonymous):

5y^2 - 125 = 0

OpenStudy (anonymous):

Standard form: 5y2 − 125 = 0

OpenStudy (anonymous):

y + 5 = 0 ⇒ y1 = −5 y − 5 = 0 ⇒ y2 = 5

OpenStudy (anonymous):

that's two ways you could do it

OpenStudy (anonymous):

do u understand?

DivineSolar (divinesolar):

5(y+5)(y−5) is the factored version.